Solve this problem through substitution:
x + y = 6
x - y = -2

Answer by Maths68(1474)

x + y = 6 ...........(1)
x - y = -2 ..........(2)
Add (1) and (2)
x + y = 6
x - y = -2
------------
2x=4
2x/2=4/2
x=2
Put the value of x in (1)
x + y = 6 ...........(1)
2+y=6
y=6-2
y=4
x,y=2,4